Basically there are three approaches to measure gdp/gnp,list and briefly explain each approach

1
Expert's answer
2022-02-06T14:22:48-0500

Solution:

GDP/GNP can be calculated in three ways: the value-added approach, the income approach (how much money is earned as income from resources used to make things), and the expenditure approach (how much is spent on stuff).


1.). The expenditures approach involves totaling all spending on final goods and services in an economy; the expenditures approach divides this spending into five categories: consumption, investment, government spending, exports, and imports: Y = C+I+G+X-M, Y = C+I+G+X-MY, equals, C, plus, I, plus, G, plus, X, minus, M.

 

2.). The income approach to GDP calculation involves adding up all of the income earned within a country's borders in a given year; the income approach includes wages, rents, interest, and profits.

 

3.). The value-added approach to GDP calculation involves adding up all of the value-added at various stages of production; for example, in the production of a $12 cake, the value-added approach counts the value of the raw ingredients that a farmer sells to the baker ($4), which a baker then combines with her capital to create a cake, which adds $8 in value.